OpenStudy (anonymous):

\[p = \sqrt{2 - p}\]

OpenStudy (anonymous):

could you better yet show me how to do them?

OpenStudy (anonymous):

p=sqrt(2-p) First, square both sides. p²=2-p Then, throw everything to one side. p²+p-2=0 Then, factor it. (p+2)(p-1)=0 p+2=0 or p-1=0 p=-2 or p=1

OpenStudy (anonymous):

this goes against every rule in algebra so far. my teacher gave every one F's because he thinks we cant learn -_-

OpenStudy (amistre64):

sqrt(4) does not equal -2

OpenStudy (anonymous):

I wasn't done yet.

OpenStudy (amistre64):

1 is good

OpenStudy (anonymous):

waiit so what is the other one?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

p=-2 is wrong, since sqrt(2-p) can't become lower than 2.

OpenStudy (amistre64):

only 1; when squareing sides, you run the risk of getting a false answer and have to check it back into the original to see if it works :)

OpenStudy (anonymous):

so then it is any number greater than 2?

OpenStudy (amistre64):

sqrt(2-p) shouldnt go below zero; below 2 is fine :)

OpenStudy (amistre64):

sqrt(2--2) = sqrt(4) = 2 -2 <> 2

OpenStudy (anonymous):

ok... this is confusing

OpenStudy (amistre64):

what is confusing about it?

OpenStudy (anonymous):

just the way my teacher told us only to do it a way i dont remember and like coding learn 2 at once you get raped... :/

OpenStudy (amistre64):

there are only two numbers whose value equals their sqrts; do you know those numbers?

OpenStudy (anonymous):

for this one?

OpenStudy (amistre64):

for any number.... does 0 = sqrt(0) ?? does 1 = sqrt(1) ??

OpenStudy (anonymous):

no sqrt for 1 0 = \[\sqrt{0}\]

OpenStudy (amistre64):

1*1 = 1 ; therefore; sqrt(1) = 1 right?

OpenStudy (anonymous):

wouldnt it equal \[\sqrt{1}\]

OpenStudy (amistre64):

1 = 1*1 = sqrt(1) = 1/1 = .....

OpenStudy (amistre64):

p = sqrt(2-p) what should we do to solve this?

OpenStudy (anonymous):

i dont even know. :/

OpenStudy (amistre64):

staring at it wont help :) Try something, even if its wrong....

OpenStudy (anonymous):

give me a minute...

OpenStudy (anonymous):

\[-2^{?} +p ^{?}\]

OpenStudy (anonymous):

ohh the question marks should be sqrd

OpenStudy (amistre64):

its a good attempt; but how do we 'do the same thing' to the other side of the equals?

OpenStudy (amistre64):

how do we 'undo' a square root?

OpenStudy (anonymous):

square it

OpenStudy (amistre64):

yes; so lets try that :) to both sides

OpenStudy (amistre64):

p^2 = sqrt(2-p)^2 p^2 = 2-p

OpenStudy (anonymous):

4 + psqrd = 0?

OpenStudy (amistre64):

do you see that steps I did?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

when you square the squarroot; sign; it disappears; that is the only thing that happens to that part....

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

sqrt(15)^2 = 15 sqrt(3)^2 = 3 sqrt(23654)^2 = 23654 see it?

OpenStudy (amistre64):

sqrt(2-p)^2 = 2-p right?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

p^2 = 2-p ; what do you wanna do to this to solve for p?

OpenStudy (anonymous):

same side?

OpenStudy (amistre64):

ok...lets put it all to the same side.... p^2 +p -2 = 2-2-p+p p^2 +p -2 = 0

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

how do we 'unFOIL' this ?

OpenStudy (anonymous):

huhm flutter if i know.

OpenStudy (amistre64):

please dont use that language here

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

i never understood the foil. my teacher writes it on the bored and says his thought erases then goes to the next prob

OpenStudy (amistre64):

FOIL is a term that is used to teach people how to multiply stuff inside of paranthesis

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

they teach it in hopes that it will mean something later on; but really its a crutch that keeps you from learning whats really happening to a problem

OpenStudy (amistre64):

The distributive property is all that we need; do you remember the distributive property?

OpenStudy (anonymous):

i ask for help bet he says a smart alic reply to me... he needs a different career

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

How do we use the distributive property then?

OpenStudy (anonymous):

times outside the ( ) by the inside of the ( )

OpenStudy (amistre64):

very good :) We can do that as many times as it takes.... watch: Do you agree that 8*3 = 24?

OpenStudy (anonymous):

yep

OpenStudy (amistre64):

then 3(5+3) should also equal 24 right? 3(5) + 3(3) = 24 15 + 9 = 24 24 = 24

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

and... (1+2)(5+3) should also equal 24..like this: (1+2)(5) + (1+2)(3) = 24 5(1) + 5(2) + 3(1) + 3(2) = 24 5 + 10 + 3 + 6 = 24 15 + 9 = 24 24 = 24 that is what they are calling FOIL..... its just using the distributive property twice

OpenStudy (anonymous):

ok so foil = complete distribution?

OpenStudy (amistre64):

yes :)

OpenStudy (anonymous):

ok.. u got a mic or something for pc? this would be easier...

OpenStudy (amistre64):

nope; using college computers..no sound no video no nothing but a keyboard

OpenStudy (anonymous):

darn... ok so please guide me through the problem im about to make a new post... itl be easier...

OpenStudy (amistre64):

ok...

OpenStudy (anonymous):

i posted it...